You run a financial service firm where you replace your employee’s computers every three years. You have 5000 employees, and each computer costs $2,500 currently-the old computers can be sold for $500 each. The new computers are generally depreciated straight-line over their three-year lives to a salvage value of $500. A computer-service firm offers to lease you the computers and replace them for you at no cost, if you will pay a leasing fee of $5 million a year (which is tax-deductible). If your tax rate is 40%, would you accept the offer?
1
Expert's answer
2012-06-21T11:06:31-0400
The amortization costs per computer are (2,500 - 500 - 500)/3 = 500 per year.&
The total amount will be 500*5000 = 2,500,000
A leasing fee of $5 million a year after taxes will be 5,000,000*(1-0,4) = 3,000,000
So we would not accept the offer, because we will save $500,000.